# LLM-ABBA: Understanding time series via symbolic approximation

### v4

@@ III-C ABBA to LLM
In the following, we write a single time series containing $n$ data points as $T$, and use $\mathcal{T}=\{T_{i}\}_{i=1}^{q}$ to denote a set of time series of cardinality $q$, associated with its corresponding symbolic representation set $\mathcal{A}=\{A_{i}\}_{i=1}^{q}$.
@@ III-C1 Fixed-point adaptive polygonal chain
In time series prediction settings, the value-based prediction is converted into a token-based prediction using STSA. However, it is very desirable to mitigate the negative effect of the preceding mistakenly predicted symbol on the subsequent time series recovery since the recovery proceeds from front to back. However, APCA and the symbolic recovery often lead to a cumulative error for symbolic prediction, that is, an incorrect replacement of a previous symbol will influence the subsequent reconstruction. A fixed-point polygonal chain trick is introduced to mitigate this issue. We still partition the time series into pieces following (1) while $p_{j}=(\text{{len}}_{j},\text{{inc}}_{j})$ is replaced with $p_{j}=(\text{{len}}_{j},t_{i_{j}})$ before normalization. We call the new approximation method fixed-point adaptive piecewise linear continuous approximation (FAPCA). The resulting tuples $p_{i}$ will be normalized and one can be recovered from the other since $\texttt{inc}_{j}=t_{i_{j}}-t_{i_{j-1}}$. Figure 4 shows that FAPCA eliminates the cumulative errors arising from the preceding mistaken symbol and improves the recovery.
@@
III-C2 Symbolizing multiple time series
Existing work on symbolic approximation focuses on converting a single time series; it can not convert another time series with consistent symbolic information (i.e., each symbol is associated with a unique symbolic center). To allow the manipulation of co-evolving time series or multiple time series, it is necessary to keep consistent symbolic information for multiple symbolic time series representations.
We illustrate a unified approach towards a consistent symbolic approximation for multiple time series.
  • Step 1: Use APCA or FAPCA to compress each time series $T_{i}$ into $P_{i}$ for $i=1,\ldots q$
  • Step 2: Compute normalized $P^{s}_{i}$ and concatenate $P^{s}_{i}$ to form $\mathcal{P}^{s}:=[P^{s}_{i}]_{i=1}^{q}$
@@ III-C2 Symbolizing multiple time series
  • Step 4: Allocate symbols to each time series (the number of symbols for $T_{i}$ is equal to $|P^{s}_{i}|$)
@@ III-C3 Symbolizing out-of-sample data
Symbolizing out-of-sample time series data with consistent symbols is essential for various time series downstream tasks, e.g., inference tasks. To symbolize $\mathcal{T}^{t}=\{T^{t}_{i}\}_{i=1}^{q^{\prime}}$, we perform the following steps:
  • Step 1: Compress each time series $T^{t}_{i}$ into $P^{t}_{i}$ for $i=1,\ldots q^{\prime}$
  • Step 2: Assign a symbol to $p\in P^{t}_{i}$ for $i=1,\ldots q^{\prime}$ following the rule of digitization
@@ III-C4 Feeding the LLM
ABBA can transform numerical time series to symbolic series and keep the internal logic chain from which LLMs can learn temporal knowledge. In other words, by ensuring the precondition that the input symbolic series inherits the polygonal chain of numerical time series and then represents this chain via symbolic series (or LLMs’ tokens) that can be recognized by LLMs, LLMs can reconstruct the embedding space without the use of any new tokens via adapting fine-tuning methods. As seen in Figure 3, the left panel is the traditional setting in terms of corresponding tasks, such as classification, regression, and prediction. The right panel is the instruction setting that contains these three tasks. Instructions (the right panel of Figure 3) only guide LLMs to understand the tasks, and this smart design is equal to the left panel (without LLMs’ Instructions) of Figure 3.
@@
    $\widehat{Y}=\operatorname{Task}\left(\mathcal{M}_{\mathrm{outp}}\right)$: If this is a classification task, $\widehat{Y}$ is a generated label. If the task is a regression or prediction task, $\widehat{Y}$ is an ABBA-transformed numerical value or sequence produced by the inverse symbolization process of ABBA:
    $\begin{cases}\widehat{Y}=\mathcal{M}_{\mathrm{outp}},&\text{Classification task},\\ \widehat{Y}=\phi^{-1}\left(\mathcal{M}_{\mathrm{outp}}\right),&\text{Regression / prediction task}\end{cases}$
@@ III-D Linguistics investigation: Zipf’s law
In nearly all corpora, the most common word appears approximately twice as frequently as the next common word; this phenomenon is explained by Zipf’s law [29]. Zipf’s law asserts that the frequencies of certain events are inversely proportional to their rank, and further, the rank-frequency distribution is an inverse power law relation.
In Figure 5, we can see unigrams generated by ABBA symbolization from 7 different time series datasets from the UCR Archive coarsely meet Zipf’s law. This showcases an appealing alignment between ABBA symbols and the native language words.
@@ IV Experiments
In this section, we study three time series tasks to validate the efficiency of ABBA in LLM. We also fine-tune three language models on the training data using QLoRA [17] with inhibition [30]. All experiments are simulated in PyTorch with a single NVIDIA A100 40GB GPU. The benefits of LLM-ABBA include (1) avoiding the need for LLMs to learn time series from scratch, and (2) only utilizing compression and decompression without the need for the training of extra embedding layers [6]. For a fair comparison, we evaluate our models on the same settings for each task. In the following, unless otherwise stated, we assume that the greedy aggregation is used for the ABBA digitization.
@@ IV-A2 Hyperparameters of LLMs
TABLE I: Hyperparameters of Classification tasks. Quant. is the model quantization process. Inhib. is the inhibition threshold in QLoRA. Embed. means to save tuned embeddings. Optims. is the optimization method. LR is the learning rate. Acc. is the accyracy rate (%).
@@
Models Quant. Tokens Metric LoRA Optim. Epochs
LR
Batch
Size
4-bits Length alpha low rank $r$ dropout inhib. Embed.
@@
Llama2-7B True 4,096 Acc. 16 16, 64, 256 0.05 0.3 Save adamw_8bit 10 5e-7 4
Mistral-7B True 4,096 Acc. 16 16, 64, 256 0.05 0.3 Save adamw_8bit 10 5e-7 4
TABLE II: Hyperparameters of Regression tasks. Quant. is the model quantization process. Inhib. is the inhibition threshold in QLoRA. Embed. means to save tuned embeddings. Optims. is the optimization method. RMSE is the root-mean-square-error.
LLM-ABBA on Regression Tasks
Models Quant. Tokens Metric LoRA Optim. Epochs
LR
Batch
Size
4-bit Length alpha low rank $r$ dropout inhib. Embed.
@@
Llama2-7B True 4,096 RMSE 16 16, 64, 256 0.05 0.3 Save adamw_8bit 10 $2\times 10^{-6}$ 4
Mistral-7B True 4,096 RMSE 16 16, 64, 256 0.05 0.3 Save adamw_8bit 10 $2\times 10^{-4}$ 4
TABLE III: Hyperparameters of Prediction tasks. Quant. is the model quantization process. Inhib. is the inhibition threshold in QLoRA. Embed. means to save tuned embeddings. Optims. is the optimization method. MAE is the mean-absolute-error, and MSE is the mean-square-error.
LLM-ABBA on Prediction Tasks
Models Quant. Tokens Metric LoRA Optim. Epochs
LR
Batch
Size
4-bit Length alpha low rank $r$ dropout inhib. Embed.
[…]
Llama2-7B True 4,096 MAE, MSE 16 16, 64, 256 0.05 0.3 Save adamw_8bit 10 $2\times 10^{-6}$ 4
Mistral-7B True 4,096 MAE, MSE 16 16, 64, 256 0.05 0.3 Save adamw_8bit 10 $2\times 10^{-6}$ 4
There are three time series analysis tasks: classification, regression, and prediction. We quantize LLMs by 4-bits using the bitsandbytes package [^2] https://github.com/bitsandbytes-foundation/bitsandbytes. In order to fine-tune LLMs, the shunting inhibition mechanism [30] is utilized during the QLoRA adapter fine-tuning progress. The modified embedding layer is also saved after fine-tuning on the corresponding task. For the classification task, the metric is accuracy rate (%). Root-mean-square-error is used as the metric for regression tasks. Mean-square-error and mean-absolute-error are used as the metrics for prediction tasks, and we also visualize the correlation coefficient of prediction tasks on ETTh1 data in terms of their seven features. We control the fine-tuning epoch and apply a small batch size on every task. The alpha of QLoRA is set to 16.
@@ IV-B Compression and Recovery
To transform the numerical time series to symbolic time series, we use tokens of LLMs as the initial dictionary of ABBA for the symbolic representation, and there are no extra tokens that will be used to represent the numerical input. ABBA shows a strong symbolic transition on time series signals (See Figure 6 and Table IV).
(a) Feaure 1.
(b) Feaure 2.
(c) Feaure 3.
(d) Feaure 4.
(e) Feaure 5.
(f) Feaure 6.
(g) Feaure 7.
Figure 6: Visualization of reconstructed input-168-predict-24 results on ETTh1 data by using ABBA symbolic approximation, where $\text{{tol}}=0.01$, $\alpha=0.01$ and $\mathrm{scl}=3$.
@@
TABLE IV: Symbolic approximation performance on ETTh1 data using ABBA. ABBA describes a time series sample by using symbolic approximation, and the number of used symbols depdnds on the complexity of the data. If the time series sample is a regular wave (for example, a sine wave), the number of used symbols is small; otherwise, ABBA needs more symbols.
ABBA Settings
@@
In this section, we observe which ABBA settings better suit time series characteristics. The default scl is set to $3$, which is used in other LLM tasks. tol and $\alpha$ are set to be the same. Table IV reports the input-168-predict-96 results when using ABBA to reconstruct ETTh1 data in terms of seven features. Setting smaller tol and $\alpha$ in ABBA can reduce the MSE and MAE scores, but more symbols or LLM tokens will be used. Under all above conditions, the correlation coefficient is 1.0.
IV-C Time Series Classification Tasks
For the classification task, we evaluate these three pretrained LLMs on UCR Time Series Archive datasets [36], EEG eye state [37], and MIT-BIH [38, 39] which have been extensively adopted for benchmarking time series classification models. We utilize cross-entropy loss for the classification training. Details of the implementation and datasets can be found in Table I. The evaluation metric is accuracy rate (%).
The UCR Archive contains 128 datasets already partitioned into train and test sets, although the ratio of the train set and test set is not always consistent [^3] The UCR Archive 2018 is available at https://www.cs.ucr.edu/~eamonn/time_series_data_2018/.. These datasets have varying numbers of labels and feature dimensions. Also, there can be uneven numbers of labels, which often results in overfitting. Therefore, classifying time series in the UCR Archive is a challenging task. Table V reports the full time series classification results on UCR2018. J1 refers to results of using k-means in the digitization process, and J2 refers to the results of using greedy aggregation (Algorithm 1) in the digitization process. We find that Algorithm 1 outperforms k-means symbolization time series transition progress in most cases.
@@ IV-C Time Series Classification Tasks
OliveOil 4 150 2.65M 66.7 46.7 12.7M 76.7 70.0 9.56M 73.3 73.3 0.3M -
PhalangesOutlinesCorrect 2 2,785 2.65M 62.2 65.4 12.7M 63.9 67.5 9.56M 62.7 67.5 0.3M -
Plane 7 1,424 2.65M 33.3 81.0 12.7M 39.0 78.1 9.56M 38.1 83.8 0.3M
PowerCons 2 2,007 2.65M 77.8 79.0 12.7M 72.8 81.1 9.56M 77.8 80.6 0.3M -
ProximalPhalanxOutlineCorrect 2 1,298 2.65M 71.5 82.8 12.7M 73.9 85.6 9.56M 72.9 83.9 0.3M -
ProximalPhalanxTW 6 1,101 2.65M 67.8 80.0 12.7M 69.8 80.0 9.56M 68.8 74.1 0.3M -
SemgHandGenderCh2 4 2,840 2.65M 49.1 54.7 12.7M 59.5 67.2 9.56M 58.3 73.3 0.3M -
@@
In the medical domain (for example, identifying the eye state using EEG signals, distinguishing abnormal ECG signals, and classifying the “normal beats”, “supraventricular ectopy beats”, “ventricular ectopy beats”, “fusion beats”, and “unclassifiable beats” of ECG signals), we report the performance of LLM-ABBA on three medical time series datasets. We set $\text{{tol}}=\alpha=0.01$. In Table VI, compared to CNN [41] on the PTB-DB data set, LLM-ABBA achieves performance almost equivalent to the SOTA. In the aspect of distinguishing MIT-BIH, CNN [41] and BiRNN [38, 42] performs the best, but LLM-ABBA slightly outperforms LSTM [43, 44].
@@
mit-bih 5 2,926 86.4 86.4 86.3 89.6 89.4 89.1 89.3 89.7 89.3 93.4 96.5 88.1
IV-D Time Series Regression Tasks

### v5

@@ III-C ABBA to LLM
In the following, we write a time series containing $n$ data points as $T_{q}$, and use $\mathcal{T}=\{T_{q}\}_{q=1}^{Q}$ to denote a set of time series of dimension $Q$, associated with its corresponding symbolic representation set $\mathcal{A}=\{A_{q}\}_{q=1}^{Q}$. If $Q=1$, $T$ is an univariate time series.
@@ III-C1 Fixed-point adaptive polygonal chain
In time series forecasting settings, the value-based forecasting is converted into a token-based forecasting using STSA. However, it is very desirable to mitigate the negative effect of the preceding mistakenly predicted symbol on the subsequent time series recovery since the recovery proceeds from front to back. However, APCA and the symbolic recovery often lead to a cumulative error for symbolic forecasting, that is, an incorrect replacement of a previous symbol will influence the subsequent reconstruction. A fixed-point polygonal chain trick is introduced to mitigate this issue. We still partition the time series into pieces following (1) while $p_{j}=(\text{{len}}_{j},\text{{inc}}_{j})$ is replaced with $p_{j}=(\text{{len}}_{j},t_{i_{j}})$ before normalization. We call the new approximation method fixed-point adaptive piecewise linear continuous approximation (FAPCA). The resulting tuples $p_{i}$ will be normalized and one can be recovered from the other since $\texttt{inc}_{j}=t_{i_{j}}-t_{i_{j-1}}$. Figure 4 shows that FAPCA eliminates the cumulative errors arising from the preceding mistaken symbol and improves the recovery.
@@
III-C2 Symbolizing multiple/multi-dimensional time series
SMTS [1] trained a tree learner to consider all attributes of multiple/multi-dimensional time series, both abbreviated as MTS, as the interactions in the space of the time index and time values can be detected from a generated high-dimensional codebook from the terminal nodes of the trees. Traditional ABBA focuses on converting a univariate time series; it can not convert MTS with consistent symbolic information (i.e., each symbol is associated with a unique symbolic center). Thus, by keeping consistent symbolic information across the dimensions of MTS, [5] proposes a joint approach to process MTS efficiently; in detail, the ABBA approach starts with concatenating the compressed pieces from separately applied adaptive piecewise polygonal chain approximation to each channel/sample and performs digitization on the merged pieces all at once, and allocates the symbol back to each channel/ sample. Based on that, the distinct symbol in each channel will natively have an implicit interaction among MTS and thus each channel/sample, which enables LLM to automatically learn it.
We illustrate a unified approach towards a consistent symbolic approximation for multiple univariate time series:
  • Step 1: Use APCA or FAPCA to compress each time series $T_{q}$ into $P_{q}$ for $q=1,\ldots Q$
  • Step 2: Compute normalized $P^{s}_{q}$ and concatenate $P^{s}_{q}$ to form $\mathcal{P}^{s}:=[P^{s}_{q}]_{q=1}^{Q}$
@@ III-C2 Symbolizing multiple/multi-dimensional time series
  • Step 4: Allocate symbols to each time series (the number of symbols for $T_{q}$ is equal to $|P^{s}_{q}|$)
@@ III-C3 Symbolizing out-of-sample data
Symbolizing out-of-sample time series data with consistent symbols is essential for various time series downstream tasks, e.g., inference tasks. To symbolize $\mathcal{T}^{\text{out}}=\{T^{\text{out}}_{q}\}_{q=1}^{Q^{\prime}}$, we perform the following steps:
  • Step 1: Compress each time series $T^{\text{out}}_{i}$ into $P^{\text{out}}_{q}$ for $q=1,\ldots Q^{\prime}$
  • Step 2: Assign a symbol to each $p\in P^{\text{out}}_{q}$ following the rule of digitization
For multivariate time series, ABBA processes them similarly to multiple univariate time series. The only additional step is to flatten each multivariate series by concatenating its values from all channels (dimensions) in sequence, creating a single one-dimensional sequence. This conversion transforms multiple multivariate time series into several univariate time series. Symbolic sequences are then generated using the same approach as for multiple univariate time series, and the symbols are mapped back to each channel (dimension).
@@ III-C4 Feeding the LLM
ABBA can transform numerical time series to symbolic series and keep the internal logic chain from which LLMs can learn temporal knowledge. In other words, by ensuring the precondition that the input symbolic series inherits the polygonal chain of numerical time series and then represents this chain via symbolic series (or LLMs’ tokens) that can be recognized by LLMs, LLMs can reconstruct the embedding space without the use of any new tokens via adapting fine-tuning methods.
For the connection to LLMs, Figure 3 shows two proposed frameworks, and only the left symmetric one is used. First, the symbolic approximation method ABBA is used to transform the time series into a symbolic series. Second, the tokenizer of large language models is utilized to tokenize the symbolic series. The quantized low-rank adaptation fine-tuning method with the shunting inhibition mechanism is applied to fine-tune the LLMs. In the classification tasks, the LLMs are loaded as a classification-following model, and the cross-entropy loss function is used. In the regression task, the bottom layer of the pre-trained LLMs is replaced by a regression layer, and the RMSE is used as the loss function to fine-tune the LLMs. In forecasting tasks, the LLMs are loaded as a generation-following model. The generated symbols will be transformed back to a numerical series by using the inverse symbolization process of ABBA. The right panel semantics is an extension application. LLMs are trained as a generation-following model, and the additional work involves incorporating more information into the instructions, such as the instruction command, task type, and domain. This instruction design is almost equal to the left panel (without LLMs’ Instructions) of Figure 3.
@@
    $\widehat{Y}=\operatorname{Task}\left(\mathcal{M}_{\mathrm{outp}}\right)$: If this is a classification task, $\widehat{Y}$ is a generated label. If the task is a regression or forecasting task, $\widehat{Y}$ is an ABBA-transformed numerical value or sequence produced by the inverse symbolization process of ABBA:
    $\begin{cases}\widehat{Y}=\mathcal{M}_{\mathrm{outp}},&\text{Classification task},\\ \widehat{Y}=\phi^{-1}\left(\mathcal{M}_{\mathrm{outp}}\right),&\text{Regression / forecasting task}\end{cases}$
The complexity of ABBA, as discussed in [12], is dominated by compression and digitization, where the compression takes linear complexity for time series of length $n$ and the digitization runtime is independent of $n$. Rather, for the compressed pieces of cardinality $k$, it uses fast aggregation which takes near-linear complexity $\mathcal{O}(k\log k+kd)$ on average and $\mathcal{O}(k^{2}d)$ in the worst case; luckily in our setting, $d=2$ is a constant since our digitization only operates on inc and len. In practice we employ a high compression rate so that $k\ll n$. To speed up the symbolization, we apply parallelism for compression which significantly decreases the runtime of the symbolization process. The training complexity of fine-tuning LLMs was described in LoRA [16]. The fine-tuning complexity of LoRA is $\mathcal{O}(n)$, but the inference complexity is $\mathcal{O}(n^{2})$. Here, $n$ is the length of the input sequence of LLMs (or the length of the pieces).
Compared to SAX’s complexity of $\mathcal{O}(n)$ and SFA’s complexity of $\mathcal{O}(n\log n)$, ABBA typically requires an additional internal data management overhead for adaptive segmentation, but it often yields superior shape preservation which benefits downstream classifiers and LLM integration.
@@ III-D Linguistics investigation: Zipf’s law
In nearly all corpora, the most common word appears approximately twice as frequently as the next common word; this phenomenon is explained by Zipf’s law [38]. Zipf’s law asserts that the frequencies of certain events are inversely proportional to their rank, and further, the rank-frequency distribution is an inverse power law relation. In Figure 5, we can see unigrams generated by ABBA symbolization from 7 different time series datasets from the UCR Archive coarsely meet Zipf’s law. This showcases an appealing alignment between ABBA symbols and the native language words.
@@ IV Experiments
In this section, we study three time series tasks to validate the efficiency of ABBA in LLM. We also fine-tune three language models on the training data using QLoRA [10] with inhibition [26]. All experiments are simulated in PyTorch with a single NVIDIA A100 40GB GPU. The benefits of LLM-ABBA include (1) avoiding the need for LLMs to learn time series from scratch, and (2) only utilizing compression and decompression without the need for the training of extra embedding layers [23]. All LLMs are quantized by 4-bits, and the used optimization method is adamw_8bit. Each data set in the corresponding tasks is fine-tuned using fewer than 10 epoches. For a fair comparison, we evaluate our models on the same settings for each task. In the following, unless otherwise stated, we assume that the sorting-based aggregation is used for the ABBA digitization.
@@ IV-A2 Hyperparameters of LLMs
TABLE I: Hyperparameters of classification tasks. Quant. is the model quantization process. Inhib. is the inhibition threshold in LoRA. Embed. means to save tuned embeddings. Optims. is the optimization method. LR is the learning rate.
@@
Models Tokens LoRA LR Batch Size
Length alpha $r$ dropout inhib. Embed.
@@
Llama-2-7B 4,096 16 16, 64, 256 0.05 0.3 Save 5e-7 4
Mistral-7B 4,096 16 16, 64, 256 0.05 0.3 Save 5e-7 4
TABLE II: Hyperparameters of regression and forecasting tasks. Quant. is the model quantization process. Inhib. is the inhibition threshold in LoRA. Embed. means to save tuned embeddings. Optims. is the optimization method.
LLM-ABBA on Regression and Forecasting Tasks
Models Tokens LoRA LR Batch Size
Length alpha $r$ dropout inhib. Embed.
@@
Llama-2-7B 4,096 16 16, 64, 256 0.05 0.3 Save 2e-6 4
Mistral-7B 4,096 16 16, 64, 256 0.05 0.3 Save 2e-6 4
There are three time series analysis tasks: classification, regression, and forecasting. We quantize LLMs by 4-bits using the bitsandbytes package [^2] https://github.com/bitsandbytes-foundation/bitsandbytes. In order to fine-tune LLMs, the shunting inhibition mechanism [26] is utilized during the QLoRA adapter fine-tuning progress. The modified embedding layer is also saved after fine-tuning on the corresponding task. For the classification task, the metric is accuracy rate (%). Root-mean-square-error is used as the metric for regression tasks. Mean-square-error and mean-absolute-error are used as the metrics for forecasting tasks, and we also visualize the correlation coefficient of forecasting tasks on ETTh1 data in terms of their seven features. We control the fine-tuning epoch and apply a small batch size on every task. The alpha of QLoRA is set to 16.
@@ IV-B Compression and recovery
To transform the numerical time series to symbolic time series, we use tokens of LLMs as the initial dictionary of ABBA for the symbolic representation, and there are no extra tokens that will be used to represent the numerical input. ABBA shows a strong symbolic transition on time series signals (See TABLE III).
@@
TABLE III: Symbolic approximation performance on ETTh1 data using ABBA. ABBA describes a time series sample by using symbolic approximation, and the number of used symbols depdnds on the complexity of the data. If the time series sample is a regular wave (for example, a sine wave), the number of used symbols is small; otherwise, ABBA needs more symbols.
Parameter Settings
@@
In this section, we observe which ABBA settings better suit time series characteristics. The default scl is set to $3$, which is used in other LLM tasks. tol and $\alpha$ are set to be the same. TABLE III reports the input-168-predict-96 results when using ABBA to reconstruct ETTh1 data in terms of seven features. Setting smaller tol and $\alpha$ in ABBA can reduce the MSE and MAE scores, but more symbols or LLM tokens will be used. Under all above conditions, the correlation coefficient is 1.0.
IV-C Robustness of LLM-ABBA on MIT-BIH data
Figure 6 presents the reconstruction of ABBA with FAPCA on MIT-BIH data that is added with Gaussian noise at signal-to-noise ratio (SNR) $\in$ $\{0dB,5dB,10dB,20dB\}$. TABLE IV shows the performance of Llama-2-7B-ABBA under different signal-to-noise ratio (SNR) cases. In the MIT-BIH data, the added Gaussian noise has a negligible impact on the performance of Llama-2-7B-ABBA.
(a) $0dB$ Gaussian noise.
(b) $5dB$ Gaussian noise.
(c) $10dB$ Gaussian noise.
(d) $20dB$ Gaussian noise.
Figure 6: Visualization of MIT-BIH with Gaussian noise at SNR $\in$ $\{0dB,5dB,10dB,20dB\}$ using ABBA.
TABLE IV: The accuracy of Llama-2-ABBA on MIT-BIH data with additive Gaussian noise at SNR $\in$ $\{0dB,5dB,10dB,20dB\}$ using Llama-2-7B.
SNR Parameter Settings Classes Symbols QLoRA
tol $\alpha$ Number Number r = 16 r = 64 r = 256
0 dB 0.01 0.01 5 4,578 89.2 89.0 88.6
5 dB 0.01 0.01 5 4,610 87.2 87.1 86.9
10 dB 0.01 0.01 5 5,391 90.0 89.7 88.7
20 dB 0.01 0.01 5 10,642 89.3 88.9 88.7
IV-D Time series classification tasks
For the classification task, we evaluate these three pretrained LLMs on UCR Time Series Archive datasets [8], EEG eye state [42], and MIT-BIH [31] which have been extensively adopted for benchmarking time series classification models. We utilize cross-entropy loss for the classification training. Details of the implementation and datasets can be found in TABLE I. The evaluation metric is accuracy rate (%).
The UCR Archive contains 128 datasets already partitioned into train and test sets, although the ratio of the train set and test set is not always consistent [^3] The UCR Archive 2018 is available at https://www.cs.ucr.edu/~eamonn/time_series_data_2018/.. These datasets have varying numbers of labels and feature dimensions. Also, there can be uneven numbers of labels, which often results in overfitting. Therefore, classifying time series in the UCR Archive is a challenging task. TABLE V reports the full time series classification results on UCR2018. J1 refers to results of using k-means in the digitization process, and J2 refers to the results of using sorting-based aggregation in the digitization process. In practice, it is observed that sorting-based aggregation outperforms k-means for time series transition progress in most cases.
@@ IV-D Time series classification tasks
In the medical domain (for example, identifying the eye state using EEG signals, distinguishing abnormal ECG signals, and classifying the “normal beats”, “supraventricular ectopy beats”, “ventricular ectopy beats”, “fusion beats”, and “unclassifiable beats” of ECG signals), we report the performance of LLM-ABBA on three medical time series datasets. We set $\text{{tol}}=\alpha=0.01$. In TABLE VI, compared to CNN [25] on the PTB-DB data set, LLM-ABBA achieves performance almost equivalent to the SOTA. In the aspect of distinguishing MIT-BIH, CNN [25] and BiRNN [43] performs the best, but LLM-ABBA slightly outperforms LSTM [41].
In this section, our primary contribution is not to achieve SOTA accuracy but to pioneer a new paradigm: enabling LLMs to natively understand time series via ABBA, using only a single consumer-grade GPU for fine-tuning. We realize that our proposed method shows only a small performance gap compared to the latest state-of-the-art results reported in Middlehurst et al. [36]. Nevertheless, it delivers highly competitive results that are within a narrow margin of the current best, thereby bridging time series classification via symbolic approximation with LLMs’ capability to capture COPs.
@@
mit-bih 5 4,578 86.4 86.4 86.3 89.6 89.4 89.1 89.3 89.7 89.3 93.4 96.5 88.1
IV-E Time series regression tasks